# Training-Free ANN-to-SNN Conversion for High-Performance Spiking Transformer

### v1

@@ Abstract
Leveraging the event-driven paradigm, Spiking Neural Networks (SNNs) offer a promising approach for constructing energy-efficient Transformer architectures. Compared to directly trained Spiking Transformers, ANN-to-SNN conversion methods bypass the high training costs. However, existing methods still suffer from notable limitations, failing to effectively handle nonlinear operations in Transformer architectures and requiring additional fine-tuning processes for pre-trained ANNs. To address these issues, we propose a high-performance and training-free ANN-to-SNN conversion framework tailored for Transformer architectures. Specifically, we introduce a Multi-basis Exponential Decay (MBE) neuron, which employs an exponential decay strategy and multi-basis encoding method to efficiently approximate various nonlinear operations. It removes the requirement for weight modifications in pre-trained ANNs. Extensive experiments across diverse tasks (CV, NLU, NLG) and mainstream Transformer architectures (ViT, RoBERTa, GPT-2) demonstrate that our method achieves near-lossless conversion accuracy with significantly lower latency. This provides a promising pathway for the efficient and scalable deployment of Spiking Transformers in real-world applications.
Index Terms:
Spiking Neural Networks, ANN-to-SNN Conversion, Spiking Transformer, Few Spikes Neuron.
I Introduction
@@
Spiking Neural Networks (SNNs) have garnered widespread attention due to their sparse spike-driven computing paradigm [1, 2]. Unlike traditional Artificial Neural Networks (ANNs) transmit information via continuous-valued activations, SNNs employ sparse binary spikes as information carriers, thereby offering superior energy efficiency for resource-limited devices [3, 4]. Recently, the SNNs community has incorporated high-performance Transformer [5] architectures into SNNs to leverage both the powerful representation capabilities of Transformer and the inherent energy efficiency of the SNNs [6, 7, 8]. These methods provide significant performance improvements and broaden the application scenarios of SNNs.
To obtain Transformer-based SNNs, two approaches exist: direct training (DT) [9, 10] and ANN-to-SNN (A2S) conversion [11, 12]. DT employs surrogate gradients [13] and backpropagation through time (BPTT), but exhibits inaccurate gradient approximation and $\mathcal{T}\times$ training overhead. In contrast, A2S converts pre-trained ANNs to SNNs, maintaining inference efficiency while avoiding training overhead. A2S methods [14, 15] achieve lossless conversion in convolutional structures by establishing an equivalence between RELU activations and firing rates [16, 17]. However, Transformer conversion remains challenging due to complex nonlinear operations.
@@
Nonlinear operations in Transformer architectures can be divided into two categories: single-variable operations (e.g., GELU, Tanh) and multi-variable operations (e.g., variable-variable floating-point (FP) multiplications, LayerNorm). These operations lead to notable performance degradation [17]. SpikeZIP-TF [18] replaces single-variable functions and applies fine-tuning to adjust pre-trained ANN weights, though at the cost of additional training overhead. STA [19] approximates all nonlinear functions using group operators, but this approach incurs increased inference latency. Therefore, it is necessary to design a high-performance conversion that balances training cost and inference efficiency.
@@
Inspired by the temporal coding mechanisms observed in biological neurons, Few-spikes (FS) neurons [20] are proposed. Unlike rate-coded approaches, it leverages spike timing and patterns to represent neuronal activation states. This characteristic enables near-lossless fitting for single-variable operations. However, vanilla FS neurons suffer from significant performance degradation in large-scale deep networks. Moreover, their design based on single-input fitting limits their ability to handle the multi-variable operations in Transformer architectures.
@@
In this paper, we propose a training-free A2S conversion framework tailored for Transformer architectures, enabling near-lossless conversion without modifying the weights of pre-trained ANNs. We begin by theoretically and empirically analyzing the key challenges associated with FS neuron-based conversion, particularly in approximating nonlinear operations. To overcome these challenges, we introduce a novel Multi-Basis Exponential Decay (MBE) neuron, which effectively approximates various nonlinear operations. Built upon MBE neurons, our method achieves both low inference latency and near-lossless Transformer conversion. The main contributions of this paper are as follows:
  1. 1.
@@
II Related work
Existing conversion paradigms can be categorized into two types based on whether extra training on the source ANNs is required: training-dependent and training-free. Training-dependent conversion typically requires additional ANNs training based on theoretical ANN-SNN equivalence. Clip-Floor-Shift activations [16] replace ReLU to better match spiking neuron behavior. A two-phase training [21] first optimizes ANN weights, then adjusts neuron thresholds, while activation range constraints and membrane potential initialization [22] help further reduce conversion error. [23] introduce a two-stage approach to handle quantization, pruning, and residual errors. [24] propose a unified framework treating spike-rate mapping as a differentiable problem. For Transformer architectures, a QANN is integrated with spatiotemporal properties of spiking neurons for lossless conversion [18]. While effective, these methods introduce computational overhead through additional source ANNs training.
Training-free conversion directly transforms pre-trained ANNs via structure and parameter reuse without fine-tuning. Early works align SNN thresholds with ANN activations via threshold balancing [11, 12], but longer timesteps are required. Signed spiking neurons [25] support dynamic vision data. In [19], training-free Transformer conversion is achieved using universal group operators and spatial rectification self-attention. SpikedAttention [26] enables spike-driven Transformer A2S via trace-based matrix multiplication and winner-take-all spike shifting, yet retains non-spiking LayerNorm. ECMT [27] reduces latency in Transformer SNNs, yet still relies on floating-point operations. While retraining is avoided, these methods often require long timesteps or retain non-spiking components, limiting SNNs’ energy efficiency.
III Preliminary & Problem Analysis
III-A Preliminary
Inspired by authentic electrophysiological characteristics of human brain neurons, FS neuron is proposed to address the inherent trade-off between spike count and accuracy in A2S [20]. By introducing learnable parameters including neuron threshold, membrane potential reset value, and spike intensity, FS neuron approximates ANN activation functions with very few spikes. Mathematically, the membrane potential of FS neuron can be computed as:
@@ III-B Problem Analysis
Figure 1: Loss landscape of FS neurons in approximating the ReLU activation function. The 3D landscape of $\mathcal{L}_{Binary}$ and $\mathcal{L}_{Random}$ from two different initialization.
Approximation Error Analysis To qualitatively analyze how the FS neuron design impacts performance, following [19], we consider errors from three sources: insufficient sampling, limited parameterization, and spiking quantization. This yields the following error bound:
@@
III-B1 Excessive Dependence on Initialization
@@ IV Method
IV-A Multi-Basis Exponential Decay Neuron
To address the limited expressiveness of FS neurons, we propose a novel MBE neuron with an exponential decay parameter update strategy and multi-basis encoding scheme. Unlike FS neurons requiring learning multiple parameters $(d[t],r[t],V_{\mathrm{th}}[t])$ per timestep, MBE neurons only learn decay rates and discrete timesteps, reducing parameter overhead and gradient instability (Details in Appendix E).
@@ IV-A Multi-Basis Exponential Decay Neuron
Approximation Error Analysis Following the same analytical framework in Theorem 1, we derive error bounds for MBE neurons, accounting for the representational capacity:
@@ IV-B Conversion Framework
The $\text{Softmax}(x_{i})=\frac{e^{x_{i}}}{\sum_{j}e^{x_{j}}}$, where $i$ ,$j$ denote the indices of elements in the sequence. Since softmax depends on all inputs, it cannot be directly implemented by a single MBE neuron. As shown in Fig. 5(b), we decompose the softmax computation into three components: exponential ($e^{x}$), reciprocal ($1/x$), and FP multiplication, each approximated by MBE neurons.
For the term $e^{x_{i}}$, we apply the change-of-base formula to decompose it into integer and decimal components as:
$e^{x_{i}}=2^{x_{i}\cdot\log_{2}e}=2^{\lfloor{x_{i}}\cdot\log_{2}e\rfloor}\cdot 2^{x_{i}-\lfloor{x_{i}}\cdot\log_{2}e\rfloor},$ (13)
where $\lfloor{\cdot}\rfloor$ denotes the floor operation. The decimal component $2^{x_{i}-\lfloor x_{i}\cdot\log_{2}e\rfloor}$ is approximated by MBE neurons, and the integer exponent $2^{\lfloor x_{i}\cdot\log_{2}e\rfloor}$ is achieved by hardware-efficient addition [28]. For the reciprocal term $1/\sum_{j}e^{x_{j}}$, we follow the IEEE 754 standard [29] to extract the exponent $E$ and mantissa $M$ through the corresponding bits of $\sum_{j}e^{x_{j}}$ (i.e., $\sum_{j}e^{x_{j}}=M\cdot 2^{E}$). The reciprocal $1/M$ is directly approximated using an MBE neuron, and the final result is obtained as $1/\sum_{j}e^{x_{j}}=2^{-E}/M$.
@@
TABLE I: ImageNet Performance Comparison.
Category Method Arch. Spike-Driven Training-Free Param[M] Timesteps ANN Acc.[%]($\Delta$)
DT DSR[30] ResNet-18 - 12 50 - 67.74
TET[31] SEW-ResNet-34 - 22 4 - 68.00
AT-SNN[32] ResNet-104 - 45 4 - 77.08
Spikingformer[33] -4-384-400E - 66 4 - 75.85
SDTv1[34] -8-768 - 66 4 - 77.07
Spikeformer[35] -7L/3×2×4 - 38 4 - 78.31
QKFormer[36] HST-10-768 - 65 4 - 84.22
SDTv3[8] E-SpikeFormer - 173 8 - 85.10
A2S QCFS [16] VGG-16 138 64 74.92 72.85(-2.07)
QFFS[37] VGG-16 138 8 73.08 73.10(+0.02)
SNM[25] ResNet-18 12 64 73.18 71.50(-1.68)
QCFS[16] ResNet-34 22 64 74.23 72.35(-1.88)
AdaFire[38] ResNet-34 22 8 75.66 72.96(-2.70)
ECL[39] ResNet-34 22 16 74.36 72.37(-1.99)
@@
ECMT[27] ViT-L/16 307 12 84.88 84.71(-0.17)
STA[19] ViT-B/32 86 256 83.60 82.79(-0.81)
SpikeZIP-TF[18] SViT-L-32Level 304 64 85.41 83.82(-1.59)
SpikedAttention[26] Swin-T(BN) 28 48 79.30 77.20(-2.10)
@@
V Experiments
V-A Experimental Setup
We conduct experiments on a GPU (RTX 4090) environment using the PyTorch framework, evaluating both pre-trained Vision Transformer (ViT) models, including ViT-Base-Patch16 (ViT-B/16) [40] and ViT-Medium-Patch16-Reg4-Gap-256 (ViT-M/16) [41], as well as CNN models, including VGG16 [42] and ResNet34 [43] on the ImageNet dataset [44]. To validate the generalizability of our method across different language domains and tasks, we adapt RoBERTa [45] for natural language understanding (NLU) tasks and employ GPT-2 [46] for natural language generation (NLG) task evaluation (Experimental and implementation details in Appendix G.1).
V-B Comparative Study
@@
We evaluate the performance of our conversion framework by applying it to convert both ViT and CNN architectures. As shown in Tab. I, ViT-B/16 and ViT-M/16 achieve conversion losses of 0.44% and 0.64% respectively—significantly lower than most existing methods. The framework attains 85.31% accuracy for ViT and 75.57% for CNN, outperforming other SNNs. While ECMT achieves competitive accuracy at shorter timesteps, it retains FP multiplication in its expectation compensation module. In contrast, our A2S framework maintains crucial spike-based property, thereby avoiding the high energy consumption typically associated FP multiplication during inference. Furthermore, the training-free nature of our approach enables conversion with minimal computational overhead, eliminating the need for extensive retraining for source ANNs. Our method successfully accomplishes A2S conversion without requiring additional training of the original network, preserves the essential spiking properties, delivers superior performance at short timesteps, and achieves optimal results across all evaluation dimensions.
@@
To evaluate our method on NLU tasks, we conduct experiments on four benchmark datasets including SST-2, SST-5, MR, and Subj, following standard experimental settings from studies [18, 47]. As shown in Table II, our method achieves competitive performance across all datasets, outperforming existing methods on both RoBERTa-Base (125M) and RoBERTa-Large (355M). On SST-2, it attains 95.98% accuracy with $T=16$, representing only 0.24% degradation from the original ANN (96.22%), while outperforming SpikeZIP-TF [18] by 2.19% and reducing timesteps by 87.5%. Similarly, on MR with RoBERTa-Base, our method achieves 89.00% accuracy at $T=16$, exceeding SpikeZIP-TF by 2.87% with 75% fewer timesteps.
Comparison on CV
We evaluate the performance of our Spike-F3 framework by applying it to convert both ViT-B/16 and ViT-M/16 architectures. The experiment results demonstrate that our Spike-F3 method achieves SOTA performance across multiple metrics. As illustrated in Tab. I, the conversion loss for the ViT-B/16 and ViT-M/16 models is measured at 0. 44% and 0. 64%, respectively, significantly lower than most existing methods, while achieving high accuracy 85.31% at T=16. Although ECMT[27] achieves good accuracy at shorter timesteps, it keeps FP multiplication in the expectation compensation module for matrix product. In contrast, converted SNNs based on Spike-$\text{F}^{3}$ maintain crucial spike-based property, thereby avoiding the high energy consumption typically associated FP multiplication during inference. Furthermore, the training-free nature of our approach enables conversion with minimal computational overhead, eliminating the need extensive retraining for source ANNs. Our method successfully accomplishes nearly lossless Transformer A2S conversion while avoiding retraining of the original ANN. It effectively preserves essential spiking neural dynamics, demonstrates superior inference accuracy under low timesteps conditions, and ultimately achieves SOTA performance across all comprehensive evaluation metrics.
TABLE II: NLU Performance Comparison
@@
DT Spikeformer[35] 110 81.55 42.02 79.38 91.80 4
SpikeBERT[48] 109 85.39 46.11 80.69 93.00 4
SpikeGPT[47] 45 80.39 37.69 69.23 88.45 50
A2S SpikeZIP-TF[18] 125 92.81 52.71 86.13 95.55 64
@@
Comparison on NLU
To assess the effectiveness of Spike-F3 in NLU tasks, we conduct experiments on four benchmark datasets including SST-2, SST-5, MR, and Subj, following standard experimental settings from previous studies [18, 47]. As shown in Tab. II, Spike-F3 achieves SOTA performance across all evaluated datasets, consistently outperforming existing methods on both RoBERTa-Base (125M) and RoBERTa-Large (355M). On SST-2, it achieves 95.98% accuracy with T=16, resulting in only 0.24% conversion loss compared to the original ANN (96.22%). Compared to SpikeZIP-TF [18], Spike-F3 improves accuracy by 2.19% and reduces timesteps by 87.5%. In the lightweight setting, Spike-F3 achieves 89.00% accuracy on MR using RoBERTa-Base with T=16, surpassing SpikeZIP-TF by 2.87% while reducing timesteps by 75%. These results confirm the effectiveness of Spike-F3 in training-free A2S for NLU tasks.
TABLE III: NLG Performance Comparison. Both Wiki-2 and Wiki-103 use perplexity. indicates lower is better.
Category Model Param [M] Wiki-2 $\downarrow$ Wiki-103 $\downarrow$ T
ANN GPT-2[46] 346 22.34 22.65 1
DT SpikeGPT[47] 216 18.01 39.75 1024
A2S Spike-F3 (Ours) 346 22.69 (+0.35) 23.41 (+0.76) 16
@@
To demonstrate the applicability of our method to NLG tasks, we evaluate its performance on WikiText-2 and WikiText-103. As shown in Tab. III, our method achieves 22.69 perplexity on WikiText-2 with only T=16 timesteps, incurring merely 0.35% conversion loss compared to GPT-2. More significantly, on WikiText-103, it attains 23.41 perplexity—a substantial 41.1% improvement over directly-trained SpikeGPT (39.75) while using dramatically fewer timesteps (T=16 vs. T=1024), closely approaching the original GPT-2 performance and demonstrating superior efficiency for long-context generation.
V-C Ablation Study
@@
TABLE IV: Performance across various timesteps and models. For Wiki-103, perplexity () is used; lower is better.
Dataset Model ANN Timestep (T)
@@
Wiki-103 GPT-2 22.65 41072 11992 33.56 23.41
We evaluate ImageNet, MR, and Wiki-103 to analyze timestep requirements for A2S conversion across vision and language tasks. As shown in Tab. IV, our method achieves about 1% conversion loss for ViT, RoBERTa, and GPT-2 at T=16, with near-optimal performance at T=12, demonstrating efficient progressive encoding. Notably, ViT-M/16 shows a 21.99% accuracy drop at T=10 vs. ViT-B/16, primarily due to its wider identity mapping range [0,62] causing amplified approximation errors under limited timesteps.
VI Energy Estimation
Unlike ANNs where energy depends on floating-point operations (FLOPs), the energy cost of SNNs is dominated by synaptic operations (SOPs). Following [49], the energy ratio between SNNs and ANNs is:
@@ VI Energy Estimation
where $E_{\text{MAC}}$ = 4.6pJ, $E_{\text{AC}}$ = 0.9pJ. For ViT-M/16, we measure the firing rates $\eta$ of nonlinear operations across all layers. Taking GELU as an example, the ANN implementation requires 70 FLOPs, while the MBE neuron achieves $\eta$=38.22% under N=4 and T=16. The synaptic operations achieve energy consumption of 13.7% according to $T*N*\eta*E_{AC}$. Other operations yield greater energy savings owing to lower MBE firing rates. Complete results and detailed firing rates are provided in Appendix G.4.
VII Conclusion

### v2

@@ Abstract
Leveraging the event-driven paradigm, Spiking Neural Networks (SNNs) offer a promising approach for energy-efficient Transformer architectures. While ANN-to-SNN conversion avoids the high training cost of directly trained Spiking Transformers, existing approaches still struggle to handle the nonlinear operations within Transformer blocks, and often require additional fine-tuning of pretrained ANNs. To address these limitations, we propose a training-free and high-performance ANN-to-SNN conversion framework tailored for Transformer architectures. Specifically, we introduce a Multi-basis Exponential Decay (MBE) neuron that combines exponential decay with a multi-basis encoding strategy to effectively approximate nonlinear operations, eliminating the need for weight modifications in pretrained ANNs. Extensive experiments across diverse tasks (CV, NLU, NLG) and mainstream Transformer architectures (ViT, RoBERTa, GPT-2) demonstrate that our method achieves near-lossless conversion accuracy with significantly lower latency. This provides a promising pathway for the efficient and scalable deployment of Spiking Transformers in real-world applications.
Introduction
@@
Spiking Neural Networks (SNNs) have garnered attention due to their sparse and spike-driven computing paradigm (Maass 1997). Unlike Artificial Neural Networks (ANNs), SNNs employ sparse binary spikes as information carriers, thereby offering superior energy efficiency for resource-limited devices (Zhang et al. 2025b; Liang et al. 2025; Wei et al. 2025). Recently, several studies (Wang et al. 2025b; Xiao et al. 2025) have integrated high-performance Transformer architectures (Vaswani et al. 2017) into SNNs to jointly exploit their expressive capacity and inherent energy efficiency. These methods substantially improve the performance of SNNs on complex tasks (Wang et al. 2025a, c; Shan et al. 2025; Cai et al. 2025).
To obtain Transformer-based SNNs, two mainstream approaches exist: Direct Training (DT) (Wu et al. 2018) and ANN-to-SNN conversion (A2S)  (Diehl et al. 2015). DT employs surrogate gradients (Neftci, Mostafa, and Zenke 2019; Wei et al. 2024; Sun et al. 2025; Zhang et al. 2025a) and backpropagation through time (BPTT) (Hochreiter and Schmidhuber 1997), making the training of large-scale SNNs feasible. However, it suffers from inaccurate gradient approximation and $\mathcal{T}\times$ training overhead. In contrast, A2S establishes an equivalence between the firing rate of Integrate-and-Fire neurons and ReLU functions (Bu et al. 2023), enabling SNNs to inherit the pretrained weights of ANNs with nearly lossless conversion in CNNs. However, complex nonlinear operations in Transformers pose substantial challenges for converting them into SNN architectures.
@@
To address this issue, prior research (Wang et al. 2023) examines nonlinear components in Transformers. According to their computational forms, these nonlinear operations can be categorized into single-variable (e.g., GELU, Tanh) and multi-variable operations (e.g., variable–variable floating-point multiplications, LayerNorm). SpikeZIP-TF (You et al. 2024) replaces single-variable functions and fine-tunes pre-trained ANN weights, incurring additional training overhead. STA (Jiang et al. 2024) approximates all nonlinear functions using group operators, which increases inference latency. Thus, there is a need for a conversion method that balances training cost with inference efficiency.
@@
Inspired by the temporal coding mechanisms observed in biological neurons, Few-spikes (FS) neurons (Stöckl and Maass 2021) are proposed. It leverages spike timing and patterns to represent neuronal activation states. This characteristic enables it to achieve near-lossless performance in small-scale CNN-based conversion (Mao et al. 2025). However, as model scale increases, FS-based conversion methods exhibit substantial performance gaps. Moreover, their single-input fitting design limits their ability to handle the multi-variable operations in Transformer architectures.
@@
In this paper, we first theoretically and empirically analyze the key challenges associated with FS–based conversion, particularly in large-scale network architectures. To overcome these challenges, we propose a training-free A2S conversion framework tailored for Transformer architectures. Specifically, we introduce a novel Multi-basis Exponential Decay (MBE) neuron that effectively approximates diverse nonlinear operations, enabling our conversion framework to achieve both low inference latency and near-lossless accuracy. The main contributions are as follows:
@@
Related Work
Existing conversion paradigms can be categorized into two types based on whether extra training on the source ANNs is required: training-dependent and training-free.
Training-dependent conversion typically requires additional training of ANN based on theoretical ANN-SNN equivalence. Clip-Floor-Shift activations (Bu et al. 2023) replace ReLU to better match spiking neuron behavior. A two-phase training (Ding et al. 2021) first optimizes ANN weights, then adjusts neuron thresholds, while activation range constraints and membrane potential initialization (Bu et al. 2022) help further reduce conversion error. (Wang et al. 2022b) introduces a two-stage approach to handle quantization, pruning, and residual errors. (Jiang et al. 2023) proposes a unified framework treating spike-rate mapping as a differentiable problem. For Transformer architectures, a QANN is integrated with the spatiotemporal properties of spiking neurons for lossless conversion (You et al. 2024). While effective, these methods introduce computational overhead through additional training of source ANNs.
Training-free conversion directly transforms pre-trained ANNs via structure and parameter reuse without fine-tuning. Early works align SNN thresholds with ANN activations via threshold balancing (Diehl et al. 2015; Rueckauer et al. 2017), but longer timesteps are required. Signed spiking neurons (Wang et al. 2022a) support dynamic vision data. In (Jiang et al. 2024), training-free Transformer conversion is achieved using universal group operators and spatial rectification self-attention. SpikedAttention (Hwang et al. 2024) enables spike-driven Transformer A2S via trace-based matrix multiplication and winner-take-all spike shifting, yet retains non-spiking LayerNorm. ECMT (Huang et al. 2024) reduces latency in Transformer SNNs, yet still relies on floating-point operations. While retraining is avoided, these methods often require long timesteps or retain non-spiking components, limiting SNNs’ energy efficiency.
Preliminary & Problem Analysis
Preliminary
Inspired by authentic electrophysiological characteristics of human brain neurons, FS neuron is proposed to address the inherent trade-off between spike count and accuracy in A2S (Stöckl and Maass 2021). By introducing learnable parameters including neuron threshold, membrane potential reset value, and spike intensity, FS neuron approximates ANN activation functions with very few spikes. Mathematically, the membrane potential of FS neuron can be computed as:
@@
Figure 1: Loss landscape of FS neurons in approximating the nonlinear activation function(i.e., ReLU). The 3D landscape of $\mathcal{L}_{Binary}$ and $\mathcal{L}_{Random}$ from two different initialization.
Approximation Error Analysis
To qualitatively analyze how the FS neuron design impacts performance, following (Jiang et al. 2024), we consider errors from three sources: insufficient sampling, limited parameterization, and spiking quantization. This yields the following error bound:
@@
Excessive Dependence on Initialization
@@
Multi-Basis Exponential Decay Neuron
To address the limitations of FS neurons, we propose a novel MBE neuron with an exponential decay parameter update strategy and multi-basis encoding scheme. Unlike FS neurons requiring learning multiple parameters $(d[t],r[t],V_{\mathrm{th}}[t])$ per timestep, MBE neurons only learn decay rates and discrete timesteps, reducing parameter overhead and gradient instability (Details in Appendix E).
@@
Approximation Error Analysis
Following the same analytical framework in Theorem 1, we derive error bounds for MBE neurons, accounting for the representational capacity:
@@
The $\text{Softmax}(x_{i})=\frac{e^{x_{i}}}{\sum_{j}e^{x_{j}}}$, where $i$ ,$j$ denote the indices of elements in the sequence. Since softmax depends on all inputs, it cannot be directly implemented by a single MBE neuron. As shown in Fig. 5(b), we decompose the softmax computation into three components: exponential ($e^{x}$), reciprocal ($1/x$), and FP multiplication, each approximated by MBE neurons. For the term $e^{x_{i}}$, we apply the change-of-base formula to decompose it into integer and decimal components as:
$e^{x_{i}}=2^{x_{i}\cdot\log_{2}e}=2^{\lfloor{x_{i}}\cdot\log_{2}e\rfloor}\cdot 2^{x_{i}\cdot\log_{2}e-\lfloor{x_{i}}\cdot\log_{2}e\rfloor},$ (13)
where $\lfloor{\cdot}\rfloor$ denotes the floor operation. The fractional part $2^{x_{i}\cdot\log_{2}e-\lfloor{x_{i}}\cdot\log_{2}e\rfloor}$ is approximated by MBE neurons, and the integer exponent $2^{\lfloor x_{i}\cdot\log_{2}e\rfloor}$ is achieved by hardware-efficient addition (Li and Gu 2023). For the reciprocal term $1/\sum_{j}e^{x_{j}}$, we follow the IEEE 754 standard (Kahan 1996) to extract the exponent $E$ and mantissa $M$ through the corresponding bits of $\sum_{j}e^{x_{j}}$ (i.e., $\sum_{j}e^{x_{j}}=M\cdot 2^{E}$). The reciprocal $1/M$ is directly approximated using an MBE neuron, and the final result is obtained as $1/\sum_{j}e^{x_{j}}=2^{-E}/M$.
@@
Category Method Arch. SD TF Param[M] Timesteps ANN Acc.[%]($\Delta$)
DT DSR(Meng et al. 2022) ResNet-18 - 12 50 - 67.74
TET(Deng et al. 2022) SEW-ResNet-34 - 22 4 - 68.00
AT-SNN(Yao et al. 2023b) ResNet-104 - 45 4 - 77.08
Spikingformer(Zhou et al. 2023) -4-384-400E - 66 4 - 75.85
SDTv1(Yao et al. 2023a) -8-768 - 66 4 - 77.07
Spikeformer(Li, Lei, and Yang 2022) -7L/3×2×4 - 38 4 - 78.31
QKFormer(Zhou et al. 2024) HST-10-768 - 65 4 - 84.22
SDTv3(Yao et al. 2025) E-SpikeFormer - 173 8 - 85.10
A2S QCFS (Bu et al. 2023) VGG-16 138 64 74.92 72.85(-2.07)
QFFS(Li, Ma, and Furber 2022) VGG-16 138 8 73.08 73.10(+0.02)
SNM(Wang et al. 2022a) ResNet-18 12 64 73.18 71.50(-1.68)
QCFS(Bu et al. 2023) ResNet-34 22 64 74.23 72.35(-1.88)
ECL(Liu et al. 2025) ResNet-34 22 16 74.36 72.37(-1.99)
AdaFire(Wang et al. 2025d) ResNet-34 22 8 75.66 72.96(-2.70)
@@
ECMT(Huang et al. 2024) ViT-L/16 307 12 84.88 84.71(-0.17)
STA(Jiang et al. 2024) ViT-B/32 88 256 83.60 82.79(-0.81)
SpikeZIP-TF(You et al. 2024) SViT-L-32Level 304 64 85.41 83.82(-1.59)
SpikedAttention(Hwang et al. 2024) Swin-T(BN) 28 48 79.30 77.20(-2.10)
@@
Table 1: ImageNet performance comparison. Here, “SD” and “TF” represent Spike-Driven and Training-Free.
Experiments
Experimental Setup
We conduct experiments on a GPU (RTX 4090) environment using the PyTorch framework, evaluating both pre-trained Vision Transformer (ViT) models, including ViT-Base-Patch16 (ViT-B/16) (Dosovitskiy et al. 2020) and ViT-Medium-Patch16-Reg4-Gap-256 (ViT-M/16) (Darcet et al. 2023), as well as CNN models, including VGG16 (Simonyan and Zisserman 2014) and ResNet34 (He et al. 2016) on the ImageNet dataset (Deng et al. 2009). To validate the generalizability of our method across different language domains and tasks, we adapt RoBERTa (Liu et al. 2019; Zhang et al. 2025c) for natural language understanding (NLU) tasks and employ GPT-2 (Radford et al. 2019) for natural language generation (NLG) task evaluation (Experimental and implementation details in Appendix G.1).
Comparative Study
@@
We evaluate the performance of our conversion framework by applying it to convert both ViT and CNN architectures. As shown in Tab. 1, ViT-B/16 and ViT-M/16 achieve conversion losses of 0.44% and 0.64% respectively—significantly lower than most existing methods. The framework attains 85.31% accuracy for ViT and 75.57% for CNN, outperforming other SNNs. While ECMT achieves competitive accuracy at shorter timesteps, it retains FP multiplication in its expectation compensation module. In contrast, our A2S framework maintains crucial spike-based property, thereby avoiding the high energy consumption typically associated FP multiplication during inference. Furthermore, the training-free nature of our approach enables conversion with minimal computational overhead, eliminating the need for extensive retraining for source ANNs. Our method successfully accomplishes A2S conversion without requiring additional training of the original network, preserves the essential spiking properties, delivers superior performance at short timesteps, and achieves optimal results across all evaluation dimensions.
@@
To evaluate our method on NLU tasks, we conduct experiments on four benchmark datasets including SST-2, SST-5, MR, and Subj, following standard experimental settings from studies (You et al. 2024; Zhu et al. 2023). As shown in Table 2, our method achieves competitive performance across all datasets, outperforming existing methods on both RoBERTa-Base (125M) and RoBERTa-Large (355M). On SST-2, it attains 95.98% accuracy with $T=16$, representing only 0.24% degradation from the original ANN (96.22%), while outperforming SpikeZIP-TF (You et al. 2024) by 2.19% and reducing timesteps by 87.5%. Similarly, on MR with RoBERTa-Base, our method achieves 89.00% accuracy at $T=16$, exceeding SpikeZIP-TF by 2.87% with 75% fewer timesteps.
@@
DT Spikeformer 110 81.55 42.02 79.38 91.80 4
SpikeBERT 109 85.39 46.11 80.69 93.00 4
SpikeGPT 45 80.39 37.69 69.23 88.45 50
A2S SpikeZIP-TF 125 92.81 52.71 86.13 95.55 64
@@
Table 2: NLU Performance Comparison
@@
To demonstrate the applicability of our method to NLG tasks, we evaluate its performance on WikiText-2 and WikiText-103. As shown in Tab. 3, our method achieves 22.69 perplexity on WikiText-2 with only T=16 timesteps, incurring merely 0.35% conversion loss compared to GPT-2. More significantly, on WikiText-103, it attains 23.41 perplexity—a substantial 41.1% improvement over directly-trained SpikeGPT (39.75) while using dramatically fewer timesteps (T=16 vs. T=1024), closely approaching the original GPT-2 performance and demonstrating superior efficiency for long-context generation.
Category Model Param Wiki-2 (↓) Wiki-103 (↓) T
ANN GPT-2 346 22.34 22.65 1
DT SpikeGPT 216 18.01 39.75 1024
A2S
Ours
346
22.69
(+0.35)
23.41
(+0.76)
16
Table 3: NLG Performance Comparison. Both Wiki-2 and Wiki-103 use perplexity, ↓ indicates that lower is better.
Ablation Study
@@
Dataset Model ANN T
@@
Wiki-103 (↓) GPT-2 22.65 41072 11992 33.56 23.41
Table 4: Performance across various timesteps and models. Wiki-103 uses perplexity, ↓ indicates that lower is better.
We evaluate ImageNet, MR, and Wiki-103 to analyze timestep requirements for A2S conversion across vision and language tasks. As shown in Tab. 4, our method achieves about 1% conversion loss for ViT, RoBERTa, and GPT-2 at T=16, with near-optimal performance at T=12, demonstrating efficient progressive encoding. Notably, the accuracy of ViT-M/16 drops to 21.99% at T=10, primarily due to its wider identity mapping range [0,62] causing amplified approximation errors under limited timesteps.
Energy Estimation
Unlike ANNs where energy depends on floating-point operations (FLOPs), the energy cost of SNNs is dominated by synaptic operations (SOPs). Following (Rathi and Roy 2020), the energy ratio between SNNs and ANNs is:
@@
where $E_{\text{MAC}}$ = 4.6pJ, $E_{\text{AC}}$ = 0.9pJ. For ViT-M/16, we measure the firing rates $\eta$ of nonlinear operations across all layers. Taking GELU as an example, the ANN implementation requires 70 FLOPs (Jiang et al. 2024), while the MBE neuron achieves $\eta$=38.22% under N=4 and T=16. The synaptic operations achieve energy consumption of 13.7% according to $T*N*\eta*E_{AC}$. Other operations yield greater energy savings owing to lower MBE firing rates. Complete results and detailed firing rates are provided in Appendix G.4.
Conclusion